OpenStudy (anonymous):

Use the substitution method to solve the linear system given. Express the answer as an ordered pair of the form (x,y) x + y = 5 2x + y = 6

OpenStudy (zzr0ck3r):

\(x+y=5\\2x+y=6\\\downarrow\\\color{blue}{x=5-y}\) plug that into the second equation \(2(\color{blue}{x})+y=6\\2(\color{blue}{5-y})+y=6\\10-2y+y=6\\10-y=6\\y=4\) plug that back into either equation to find \(x\) \(x=5-y=5-4=1\) so our solution is \((1,4)\).